Define variables, state the equation. The width of a rectangular room is four feet less than its length. If the area of the room is 252 square feet, what are the dimensions?

Answer by Cromlix(4381)

Hi there,
Make length = x
Width = x - 4
Area of rectangle
= width x length
252 = (x - 4) * x (* times)
x^2 - 4x = 252
x^2 - 4x - 252 = 0
(x - 18)(x + 14) = 0
x + 14 = 0
x = -14 (Disregard as -ve value)
x - 18 = 0
x = 18
Length = 18 feet
Width = 14 feet.
Hope this helps :-)